OpenStudy (anonymous):

What is the x-intercept of the line given by the equation x-4=20?

OpenStudy (anonymous):

whats 20+4?

OpenStudy (anonymous):

24 how did you get that?

OpenStudy (anonymous):

it's asking for the x-intercept x-4=20 you need to get x by itself so you would add 4 to 24 so it's x=24 x=x-intercept

OpenStudy (anonymous):

x-4y=20

OpenStudy (anonymous):

-4y=-x+20 y=1/4x-5 y=mx+b m being slope b being y intercept whats slope and y intercept?

OpenStudy (anonymous):

What is the x-intercept of the line given by the equation x-4y=20

OpenStudy (anonymous):

i put the equation in y=mx+b form thats slope m= slope of the line b=y-intercept

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 4}x-5\]